OpenStudy (anonymous):

I need someone to work this out completely and show all the steps using the formula, H(t)=-16t^2+vt+s. A juggler is performing an act by juggling several balls. The juggler throws the balls up at an initial height of 3.9 feet with a speed of 14.8 feet per second. If the juggler did not catch a ball, about how long will it take the ball to hit the floor?

OpenStudy (cwrw238):

H(T) is the height of ball at time t when ball hits ground tho is 0 s is initial height = 3.9 ft v = 14.8 ft/s - the initial speed so we have 0 = 16t^2 + 14.8t + 3.9 solving for t will give you the time you require

OpenStudy (cwrw238):

thats a quadratic equation you can solve it using the quadratic formula - ar you familiar with this?

OpenStudy (cwrw238):

for the equation ax^2 + bx + c = 0 x = [-b +- sqrt(b^2-4ac)] / 2a

OpenStudy (anonymous):

Wait so you just plug it in and solve for t? Because other people were saying you can do it another way and it took alot less time.

OpenStudy (cwrw238):

here a = 16, b = 14.8 and c = 3.9 x = [-14.8 +/- sqrt(14.8^2 - 4*16*3.9)\ / 32 hmm - somethins wrong - this won't work out as we have square root of a negative number r u sure that formula is correct?

OpenStudy (cwrw238):

oh sorry - its ny mistake the formula is -16t^2 not 16t^2

OpenStudy (cwrw238):

its 0 = -16t^2 + 14.8t + 3.9 thats same as 16t^2 - 14.8t - 3.9 = 0 can you try and solve this by the method i've shown?

OpenStudy (cwrw238):

t = 14.8 + sqrt(14.8^2 - 4*16*-3.9) / 32 = (14.8 + 21.65 )/ 32 = 2.28 seconds theres also negative root which we can ignore the answer is 2.28 seconds

OpenStudy (anonymous):

I am so lost.. I never got taught this correctly I tried to teach myself and that hasn't gotten me very far.

OpenStudy (cwrw238):

yes - quadratic equations are a tricky when you come across them first time there must be some video tutorial about them online try googling quadratic equations

OpenStudy (cwrw238):

try khan academy

OpenStudy (cwrw238):

search for 'khan academy algebra' should do it